Rocky Mountain Tire Center sells 11,000 go-cart tires per year. The ordering cost for each order is $40, and the holding cost is 50% of the purchace price of the tires per year. The purchase price is $25 per tire if fewer than 200 tires are ordered, $18 per tire is 200 or more - but fewer than 8000 - tires are orderd and $13 per tire if 8,000 or more tires are ordered.

a. How many tires should Rocky Mountains tire order each time it time it places an order?

b. What is the total cost of this policy?
Answer #1

Given that, Yearly demand = 11000 Units

Ordering cost = $40 per Order

Let the pyrchase price used = $25 per unit

Holding cost = $50% of 25 = $12.5 per unit per year

We know that,

Economic order quantity = EOQ = Sq. Root [(2 x Annual demand x Ordering cost) / Holding cost]

= Sq. Root [(2 x 11000 x 40) / 12.5]

= 265.33 Units

Total annual orders = Annual demand / EOQ = 11000 / 265.33 = 41.46 Orders

Annual ordering cost = 41.46 orders x $40 per order =$1658.31

In EOQ, Annual ordering cost = Annual holding cost = $1658.31

Total purchase cost = 11000 Units x 18 per unit = $198,000

Total cost = Total purchase cost + Annual ordering cost + Annual holding cost

= 198000 + 1658.31 + 1658.31

= $201316.62